x Give Up? t Score: 10 of 19 > The Acmeville Metropolitan Bus Service currently charges $0.88 for an all-day ticket and is used by an average of 513 riders a day. The bus company is not earning a profit, but according to their contract with the city, they cannot cut the number of buses on the road. They must, therefore, find a way to increase revenues The bus company is considering increasing the ticket price to $0.99. The marketing departments studies indicate this price increase would reduce usage to 363 riders per day Calculate the price clasticity of demand using the midpoint method for bus tickets to determine if the bus company should increase price or decrease price to increase revenues. Enter your answer as an absolute value and round it to two places after the decimal. price elasticity of demand:-2.88 Determine if demand is elastic or inelastic and what this implies regarding how ticket prices affect revenue. O Demand is elastic, so decreasing ticket prices will increase revenue O Demand is inelastic, so decreasing ticket prices will increase revenue. O Demand is inelastic, so increasing ticket prices will increase re O Demand is elastic, so increasing ticket prices will increase revenue. DOLL
Answer #1

P1 = 0.88, Q1 = 513, P2 = 0.99 and Q2 = 363

Elasticity of demand = % change in quantity demanded / % change in price

= [(Q2 - Q1)/{(Q1 + Q2)/2}] X 100 / [(P2 - P1)/{(P1 + P2)/2}] X 100

= (P1 + P2)/(Q1 + Q2) x (Q2 - Q1)/(P2 - P1)

= (0.88 + 0.99)/(513 + 363) x (363 - 513)/(0.99 - 0.88)

= 1.87/876 x - 150/0.11 = - 2.91

Demand is elastic because Ed is greater than 1.

Increase in price decreases total revenue while decrease in price increases total revenue.

Demand is elastic, so decreasing ticket prices will increase total revenue.